### v1

# An Effective Way to Determine the Separability of Quantum State

Abstract
We propose in this work a practical approach, by virtue of correlation matrices of the generic observables, to study the long lasting tough issue of quantum separability. Some general separability conditions are set up through constructing a measurement-induced Bloch space. In essence, these conditions are established due to the self constraint in the space of quantum states. The new approach can not only reproduce many of the prevailing entanglement criteria, but also lead to even stronger results and manifest superiority for some bound entangled states. Moreover, as a by product, the new criteria are found directly transformable to the entanglement witness operators.
Introduction
Quantum entanglement captures the most enigmatic and deepest insights into quantum theory, and has deep connections with other branches of physics [1]. In quantum information theory, how to distinguish an entangled state from a separable state is a fundamental and significant problem, say the so-called separability problem [2, 3]. It is well-known that the separability problem tends to be a NP-hard issue with the increase of dimension, even for the bipartite system [4], to develop a practical and experimental friendly separability criterion is pivotal in the establishment of a complete entanglement theory [2, 3]. In past decades, numerous methods have been put forward to attack it, including positive map theory and entanglement witness [5, 6, 7], the computable cross norm or realignment (CCNR) criterion [8, 4, 3], correlation matrix method [11, 12, 13, 5], local uncertainty relation (LUR) criteria [15, 16, 17, 18, 19, 20], covariance matrix [21, 22], quantum Fisher information (QFI) method [23, 24, 25], moment method [26, 27, 28, 29], etc.
In this Letter, we develop a framework to derive separability criterion for finite-dimensional systems in terms of the correlation matrix. To realize it, a measurement-induced, generalized Bloch space for the arbitrary observables is introduced, and by which a generic separability criterion is formulated. The key point here is that the correlation matrix of the separable state is limited by the size of the generalized Bloch space. The powerfulness of this approach can be realized by noticing that in which the de Vicente’s correlation matrix criterion [11], CCNR criterion [4, 3], the correlation matrix criterion based on symmetric informationally complete positive operator-valued measures (SIC-POVM) [5] etc. can all be reproduced transparently, and even stronger criterion may be obtained. Moreover, the new scenario also provides a concise method to construct an entanglement witness operator via generic observables. Examples are provided to show the availability of bound entangled state detection.
Bloch representation of observables and density matrix
The notations to be employed in the discussion go as follows. The observable is denoted by a capital letter, such as $A$, and a boldface capital letter $\boldsymbol{A}$ implies a measurement vector containing certain observables, i.e. $\boldsymbol{A}=(A_{1},A_{2},\cdots)$. The density matrix $\rho$ of a $d$-level quantum system is formulated in Bloch representation, to wit [30, 31, 32, 33, 35, 34]
$\rho=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{r}\cdot\vec{\pi}\;.$ (1)
Here, $\vec{\pi}$ are $d^{2}-1$ generators of $\mathfrak{su}(d)$ Li algebra. Due to the orthogonality relation of the generators $\mathrm{Tr}[\pi_{\mu}\pi_{\nu}]=2\delta_{\mu\nu}$, we have $r_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}]$. The set of the Bloch vectors $\vec{r}$ constitutes the so-called Bloch space. The positive semidefiniteness of density matrix constrains the size of Bloch space, i.e. $|\vec{r}|\leq\sqrt{2(d-1)/d}$. Similarly, the Bloch representation of an observable $A_{\mu}$ reads
$A_{\mu}=$ $\frac{t_{\mu}}{d}\mathds{1}+\vec{a}_{\mu}\cdot\vec{\pi}$
[…]
where $t_{\mu}=\mathrm{Tr}[A_{\mu}]$ and $a_{\mu\nu}=(\vec{a}_{\mu})_{\nu}=\mathrm{Tr}[A_{\mu}\pi_{\nu}]/2$. And then, a measurement vector $\boldsymbol{A}$ can be parameterized as a real matrix $M$ in Bloch representation
$\boldsymbol{A}\rightarrow M_{A}\equiv\left(\begin{matrix}\frac{t_{1}}{\sqrt{2d}}&\frac{t_{2}}{\sqrt{2d}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;.$ (5)
Note, here $\vec{a}_{\mu}$ are column vectors.
Correlation matrix of separable state and measurement-induced Bloch space
Quantum entanglement exhibits inherent non-classical correlation in many-body systems, even far apart from each other, which was initially recognized by Einstein, Podolsky and Rosen [36] and Schrödinger [37]. Based on correlation function [38] Bell inequality provides a practical way for the first time to test the entanglement. In the family of Bell inequality, the Clauser-Horne-Shimony-Holt (CHSH) inequality [39]
$\left|\braket{A\otimes B}-\braket{A\otimes B^{\prime}}+\braket{A^{\prime}\otimes B}+\braket{A^{\prime}\otimes B^{\prime}}\right|\leq 2\;$ (6)
might be the most prominent one, which tends to be experiment-operational friendly. Here, $A(A^{\prime})$ and $B(B^{\prime})$ are pairs of dichotomic observables with eigenvalues of ±$\pm 1$ measured by the parties $A$ and $B$, respectively. The violation of CHSH inequality and the predictions of quantum theory were first confirmed experimentally by [40, 41, 42], then later more convincingly by some loophole-free tests [43, 44, 45, 46]. The violation of CHSH inequality implies the entanglement and exclusion of local hidden variable scenario for the description of the correlation system. However, in reverse, Bell inequalities fail to recognize all the entangled states [47]. Hence, to explore and exploit the nature of entanglement, it is essential to find even stronger criteria for the determination of separable state.
To this end, we define the following correlation matrix entry for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$
$\mathcal{C}_{\mu\nu}\equiv\braket{A_{\mu}\otimes B_{\nu}}\;.$ (7)
Generally speaking, a state is separable iff it is a convex combination of product states, i.e. $\rho_{\mathrm{sep}}=\sum_{i}q_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$, where $\{\rho_{i}^{A}\}$ and $\{\rho_{i}^{B}\}$ are two pure subsystems. And, hence the corresponding correlation matrix for a separable state can write
$\mathcal{C}=\sum_{i}q_{i}\left(\begin{matrix}\braket{A_{1}}_{i}\\ \braket{A_{2}}_{i}\\ \vdots\end{matrix}\right)\left(\begin{matrix}\braket{B_{1}}_{i},\braket{B_{2}}_{i},\cdots\end{matrix}\right)\;,$ (12)
where $\braket{A_{\mu}}_{i}=\mathrm{Tr}[\rho_{i}^{A}A_{\mu}]$ and $\braket{B_{\nu}}_{i}=\mathrm{Tr}[\rho_{i}^{B}B_{\nu}]$. Eq. 12 inspires us to define a measurement-induced Bloch vector(MIBV) $\vec{\alpha}$ for measurements $\boldsymbol{A}$, of which
$\alpha_{\mu}\equiv\mathrm{Tr}[\rho\bar{A}_{\mu}]=\vec{r}\cdot\vec{a}_{\mu}\;$ (13)
with $\bar{A}_{\mu}=\vec{a}_{\mu}\cdot\vec{\pi}$ denoting the traceless part of observable $A_{\mu}$. And, various MIBV $\vec{\alpha}$ span the corresponding measurement-induced Bloch space, signified as $\mathcal{B}(\boldsymbol{A})$, which has been applied to the study of certainty and uncertainty relations [1]. The correlation matrix of a separable state can be reformulated by MIBVs, like
$\mathcal{C}=\sum_{i}q_{i}(\vec{t}^{A}/d_{A}+\vec{\alpha}_{i})(\vec{t}^{B}/d_{B}+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (14)
Here, $\vec{t}^{A}$ and $\vec{t}^{B}$ are column vectors consisting of traces of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, respectively; $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are MIBVs of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$; $\mathrm{T}$ denotes the transpose of a given matrix.
The separability criteria
A family of sufficient conditions for entanglement of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$ goes as follows, with proof given in the Supplement:
Theorem 1.
Let $\mathcal{C}$ be a correlation matrix for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, if the corresponding state $\rho$ is separable, then
$\|\mathcal{C}\|_{\operatorname{tr}}\leq\kappa_{A}\kappa_{B}\;.$ (15)
Otherwise, $\rho$ will be an entangled state. Here, $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$ is the trace norm of matrix $X$; $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ with $\lambda_{\max}(X)$ being the maximal eigenvalue of matrix $X$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, and similarly for $\kappa_{B}$.
Next, we illustrate how Theorem 1 works. For a bipartite system, we have the following Bloch representation
$\rho$ $=\frac{1}{4}\sum_{\mu,\nu}\chi_{\mu\nu}\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B}$ (16)
$=\frac{1}{d_{A}d_{B}}\mathds{1}\otimes\mathds{1}+\frac{1}{2d_{B}}\vec{a}\cdot\vec{\pi}^{A}\otimes\mathds{1}+\frac{1}{2d_{A}}\mathds{1}\otimes\vec{b}\cdot\vec{\pi}^{B}+\frac{1}{4}\sum_{\mu,\nu}\mathcal{T}_{\mu\nu}\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}\,.$ (17)
Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$; $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d^{2}-1}$ are generators of $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,\ B$; the coefficient matrix $\chi_{\mu\nu}=\mathrm{Tr}[\rho(\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B})]$; $a_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\mathds{1}]$; $b_{\mu}=\mathrm{Tr}[\rho\mathds{1}\otimes\pi_{\mu}^{B}]$; $\mathcal{T}_{\mu\nu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}]$; and $\chi_{00}=\frac{2}{\sqrt{d_{A}d_{B}}}$ since $\mathrm{Tr}[\rho]=1$. The correlation matrix can then be reexpressed as
$\mathcal{C}=M_{A}^{\mathrm{T}}\chi M_{B}\;$ (18)
with $M_{A}$ and $M_{B}$ being the Bloch representations of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ defined in Eq. 5. Equipped with this, one may notice that the traceless part of the measurement $\boldsymbol{A}$ can be viewed as the vertex of a polygon or polyhedron in $(d^{2}-1)$-dimensional space, which is enlightening for further processing of Theorem 1.
Consider the simplest situation $t_{\mu}^{A}=0$, and $\vec{a}_{1}=\vec{0}$, $\vec{a}_{\mu}=(0,\cdots,0,1,0,\cdots,0)^{\mathrm{T}}$, $\mu=2,\cdots,d_{A}^{2}$, where all components of $\vec{a}_{\mu}$ equal to 00, except the $\mu$th equals to $1$. It is easy to find $\lambda_{\max}(\Omega_{A})=2$ (similarly for $B$) and the separability condition
$\|\mathcal{C}\|_{\mathrm{tr}}=\|\mathcal{T}\|_{\mathrm{tr}}\leq\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}\;.$ (19)
Here, $\mathcal{T}=\mathrm{Tr}[\rho(\pi_{\mu}\otimes\pi_{\nu})]$. Note, Eq.(19) simply reproduces the de Vicente’s correlation matrix criterion [11].
Moreover, if we let $t_{1}^{A}=\sqrt{2d}h_{A},t_{1}^{B}=\sqrt{2d}h_{A},t_{\mu}^{A}=t_{\mu}^{B}=0,\mu\neq 1$, and take $\{\vec{a}_{\mu}\}$ and $\{\vec{b}_{\mu}\}$ the same as in above, we obtain
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{\frac{2(d_{A}-1+h_{A}^{2})}{d_{B}}}\sqrt{\frac{2(d_{B}-1+h_{B}^{2})}{d_{B}}}\;,$ (20)
which is exactly the Sarbicki et al.’s criterion [6]. Whereas, if $\boldsymbol{A}$, and similarly for $\boldsymbol{B}$, has the same trace and its traceless part corresponds to a regular simplex, viz.
$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (23)
where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e. $\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\}$ for $\mu,\nu=1,\cdots,d_{A}^{2}$, a family of separability criteria can be readily obtained (the proof of separability criteria and construction of regular simplex vertex coordinates are given in Supplements):
Observation 1.
Given measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$ and traceless parts correspond respectively to a regular simplex. The correlation matrix of separable state $\rho$ then satisfy
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;,$ (24)
otherwise, $\rho$ is entangled. Here, parameters $t_{A}$ and $t_{B}$ are arbitrary real numbers.
The 1 represents a family of separability conditions for different $t_{A}$ and $t_{B}$, many of existing criteria appear to be its particular cases. When $t_{X}$ takes $\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $\sqrt{\frac{2d_{X}}{d_{X}-1}}$ ($X=A,B$), it gives out the well-known CCNR [50, 4] and entanglement criterion via SIC-POVM (ESIC) criteria [5], respectively. To exhibit more, following we compare 1 with some symbolic criteria under certain quantum states, the Horodecki’s bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$, unextendible product bases (UPB) bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and chessboard states $\rho_{\mathrm{CB}}$. The exact forms of these quantum states are shown in the Supplemental Material.
In Fig. 1 we show and compare 1, CCNR and ESIC criteria under three typical classes of bound entangled states, where larger $t$ witnesses higher robustness against white noise and hence at which larger portion of entangled state may be detected. Fig. 1(a) exhibits Horodecki’s bound entangled states with white noise and Fig. 1(b) explains the detection results for UPB state with white noise (peru solid line) and $50000$ randomly generated chessboard state (blue dashed line). Note, as shown in the figure, 1 reproduces the CCNR and ESIC criteria as two special cases when $t=\sqrt{3}/2$ and $t=\sqrt{3}$, respectively.
(a) Horodecki’s bound entangled states with white noise
(b) UPB and chessboard states
Figure 1: Entanglement detection criteria. (a) Horodecki’s $3\times 3$ bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$. (b) UPB bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and $50000$ randomly generated chessboard states $\rho_{\mathrm{CB}}$. Here, we simply take $t_{A}=t_{B}=t$. The larger $t$ detects the more entangled states as shown in (a) and (b), in which $t=\sqrt{3}/2$ and $t=\sqrt{3}$ correspond to CCNR and ESIC criteria, respectively. Diagrams indicate explicitly that the criterion given in Observation 1 outperform CCNR and ESIC criteria under the three typical classes of bound entangled states.
Encouraged by the results in Fig.1, one may naturally ask whether larger $t$ means a more powerful criterion, or, specifically, whether the ESIC criterion is stronger than the CCNR or not, which was conjectured to be true in Ref. [5]. Unfortunately, Eq. 45 can not tell, since the left hand side (lhs) of it contains as well parameters $t_{A}$ and $t_{B}$. Nevertheless, we find this question can be readily addressed under the local filtering transformation (LFT), which preserves the separability of a given state, i.e. $\rho\xmapsto{}\tilde{\rho}=(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}/\mathrm{Tr}[(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}]$ with $F_{A/B}$ the arbitrary invertible matrices [51]. LFT transforms the full local rank state into a normal form with maximally mixed subsystems [51, 52]
$\tilde{\rho}=\frac{1}{d_{A}d_{B}}\mathds{1}\otimes\mathds{1}+\frac{1}{4}\sum_{\mu,\nu}\widetilde{\mathcal{T}}_{\mu\nu}\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}\;.$ (25)
By dint of quantum states formulated in form of Eq. 25 the bipartite separability problem [52] will be solved. Concretely, we have the following theorem (see Supplemental Material for proof):
Theorem 2.
Let measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ satisfy $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$. If a quantum state $\rho$ is separable, one may find the following relation
$\mathrm{Tr}\left[\sqrt{\mathcal{A}\widetilde{\mathcal{T}}\mathcal{B}\widetilde{\mathcal{T}}^{\mathrm{T}}}\right]\leq\kappa_{A}\kappa_{B}-\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}\;,$ (26)
where $\mathcal{A}=\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}$ and $\mathcal{B}=\sum_{\mu}\vec{b}_{\mu}\vec{b}_{\mu}^{\mathrm{T}}$.
As a result of LFT, Theorem 2 formulates a family of even stronger separability criteria than Theorem 1. If $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, then we readily have $\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}$ and the following Observation:
Observation 2.
Given measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$, and traceless parts correspond respectively to a regular simplex. If a quantum state $\rho$ is separable, then in its normal form $\tilde{\rho}$
$\|\widetilde{\mathcal{T}}\|_{\mathrm{tr}}\leq\kappa(t_{A},t_{B})\;.$ (27)
Here,
$\kappa(t_{A},t_{B})=\frac{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}{d_{A}d_{B}}\left(\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}-|t_{A}t_{B}|\right)\;.$ (28)
Now that the lhs of Eq. 27 is independent of parameters $t_{A},t_{B}$, one can then strictly compare it with other criteria for different $t_{A}$ and $t_{B}$. Notice that the minimal value of $\kappa(t_{A},t_{B})$, i.e. $\min_{\{t_{A},t_{B}\}}\kappa(t_{A},t_{B})=2\sqrt{(d_{A}-1)(d_{B}-1)/d_{A}d_{B}}$, corresponds to the strongest constraint on separability, which reproduces the Proposition 6 of Ref. [21] when $d_{A}=d_{B}$. Furthermore, the CCNR and ESIC criteria will reexhibit respectively while we take $\kappa(t_{A},t_{B})=2-2/\sqrt{d_{A}d_{B}}$ and $4-2\sqrt{(d_{A}+1)(d_{B}+1)/d_{A}d_{B}}$, which also implies that in normal form the ESIC criterion is stronger than CCNR and they coincide as $d_{A}=d_{B}$.
Relations to entanglement witness
An entanglement witness is a Hermitian operator which has nonnegative expectation values for all separable states and negative one for some entangled states [7, 53]. Entanglement witness provides an experimental friendly tool to analyse entanglement [3]. To this end, we reformulate the Theorem 1 in terms of entanglement witness. Given a correlation matrix $\mathcal{C}$ for measurement vectors $\boldsymbol{A},\boldsymbol{B}$, we readily find
$\|\mathcal{C}\|_{\mathrm{tr}}=\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\;.$ (29)
Here, $A^{\prime}_{\mu}=\sum_{\nu}P_{\nu\mu}A_{\nu}$, $B^{\prime}_{\mu}=\sum_{\nu}Q_{\nu\mu}B_{\nu}$ and $P,Q$ achieve a singular value decomposition of $\mathcal{C}$, i.e. $\mathcal{C}=P\Sigma Q^{\mathrm{T}}$, where $P,Q$ are orthogonal matrices, $\Sigma=\operatorname{diag}\{\sigma_{1},\cdots,\sigma_{r},0,\cdots,0\}$ with singular values $\sigma_{k}$ and $\operatorname{rank}(\mathcal{C})=r$ [54]. Thus, one may immediately get an inequality of correlation function similar to CHSH inequality Eq. 6 for separable state
$\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\leq\kappa\;,$ (30)
where $\kappa=\kappa_{A}\kappa_{B}$. Obviously, Eq. 30 can be reformulated as the following entanglement witness
$\mathcal{W}=\kappa\mathds{1}\otimes\mathds{1}-\sum_{\mu=1}^{r}A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}\;.$ (31)
This means that based upon the extended Bloch space, entanglement witness will be simply generated for generic observables.
Further extensions
Through the construction of a measurement-induced Bloch space, we have established a general condition for separability which arises from the constraint of the length of the MIBVs. However, not only the moduli of MIBVs but also their directions suffice a constraint. As an illustration, we consider the measurement vector $\boldsymbol{A}=\boldsymbol{B}=\vec{\pi}$, in which case the measurement-induced Bloch space reduces to the familiar Bloch space. Then the angel $\theta$ between any two Bloch vectors of pure states satisfies $-\frac{1}{d-1}\leq\cos\theta\leq 1$ [55, 35]. Given a $d\times d$ bipartite state, if it is separable, i.e. $\mathcal{T}=\sum_{i}p_{i}\vec{r}_{i}\vec{s}_{i}^{\mathrm{T}}$ with $\vec{r}_{i}$ and $\vec{s}_{i}$ being the Bloch vectors, we find the criterion
$-\frac{2}{d}\leq\mathrm{Tr}[\mathcal{T}]\leq\frac{2(d-1)}{d}$ (32)
exists, otherwise the entanglement survives. We notice $\mathrm{Tr}[\mathcal{T}]=\sum_{\mu=1}^{d^{2}-1}\braket{\pi_{\mu}\otimes\pi_{\mu}}$, so Eq. 32 yields the entanglement witness $\mathcal{W}=\frac{2}{d}\mathds{1}\otimes\mathds{1}+\sum_{\mu}\pi_{\mu}\otimes\pi_{\mu}$. Considering arbitrary dimensional Werner states [47, 56]
$\rho_{\mathrm{W}}=\frac{1}{d^{2}}\mathds{1}\otimes\mathds{1}+\frac{1}{4}\sum_{\mu=1}^{d^{2}-1}\frac{2(d\phi-1)}{d(d^{2}-1)}\pi_{\mu}\otimes\pi_{\mu}\;,\;\forall\phi\in[-1,1]\;,$
which is separable if and only if $\phi\geq 0$ [47]. An simple calculation shows that Eq. 32 is a necessary and sufficient condition taking effect in full parameter space $\phi\in[-1,0)$ , while CCNR and ESIC criteria are valid merely in $\phi\in[-1,-\frac{d-2}{d}]$. With the growth of dimensionality, the entanglement detectable intervals of CCNR and ESIC criteria tend to none. This example tells that we can establish more powerful separability condition than existing ones by only employing the inherent constraints in Bloch space. On this account, we are tempting to conclude that the inherent constraints in Bloch space may address a lot of the separability issues.
Concluding remarks
In the framework of measurement-induced Bloch space, by dint of a novel scheme we obtain a family of stronger separability criteria for entangled system. These criteria encompass many of the well-known existing results, meanwhile explicitly show why they work, i.e. the correlation matrix of a separable state is constrained by the measurement-induced Bloch space. Moreover, the new criteria can be transformed into entanglement witness operators, which makes the construction of entanglement witness from generic observables even transparent. Finally, It is noteworthy that the method developed in this work is in principle extendable to the multipartite systems by introducing the corresponding multipartite correlation tensors, though some tedious work needs to be done.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China(NSFC) under the Grants 11975236, 12235008, the Fundamental Research Funds for the Central Universities and by the University of Chinese Academy of Sciences.
Supplemental Material
Here we elaborate and derive in detail the main results discussed in the main text.
The proof of Theorem 1
Given arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, the correlation matrices of the separable states can be expressed as
$\mathcal{C}=\sum_{i}p_{i}(\vec{t}\,^{A}/d_{A}+\vec{\alpha}_{i})(\vec{t}\,^{B}/d_{B}+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (33)
Here, $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are the generalized Bloch vectors associated with $\boldsymbol{A}$ and $\boldsymbol{B}$ as defined in the main text. Employing the convexity property of trace norm $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$, we have
$\|\mathcal{C}\|_{\mathrm{tr}}$ $\leq\sum_{i}p_{i}|\vec{t}\,^{A}/d_{A}+\vec{\alpha}_{i}||\vec{t}\,^{B}/d_{B}+\vec{\beta}_{i}|$ (34)
$=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\vec{\alpha}_{i}^{2}}\sqrt{(\vec{t}\,^{B})^{2}/d_{B}^{2}+\vec{\beta}_{i}^{2}}\;.$ (35)
Here, we make use of $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$ and ||$\cdot$|| ⋅ | denotes Euclidean norm of a real vector. So, the correlation matrices of the separable states are limited by the size of the measurement-induced Bloch space (MIBS). Next, we prove that the MIBS is bounded and its boundary can be obtained simply.
Given a measurement vector $\boldsymbol{A}$, quantum states can be transformed into the measurement-induced Bloch vector (MIBVs) which constitute the MIBS. It is found that the MIBVs satisfy the following equation [1]
$\vec{\alpha}^{\mathrm{T}}\Omega^{-}\vec{\alpha}=\mathrm{Tr}[\rho^{2}]-\frac{1}{d}\;.$ (36)
Here, $\Omega_{\mu\nu}\equiv\mathrm{Tr}[\bar{A}_{\mu}\bar{A}_{\nu}]=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$ and $\Omega^{-}$ denotes the Moore-Penrose inverse of $\Omega$. Eq. 36 is an ellipsoid equation, which characterizes the MIBS, that is,
$\mathcal{B}(\boldsymbol{A})=\left\{\vec{\alpha}\Big{|}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}^{\mathrm{T}}\Omega^{-}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}\leq 1\right\}\;.$ (37)
Matrix $\Omega$ completely depicts the properties of the ellipsoid defined by Eq. 37, such as the lengths of the semi-axes are given by $\sqrt{\frac{(d\mathrm{Tr}[\rho^{2}]-1)\lambda_{i}}{d}}$, with $\lambda_{i}$ being the eigenvalues of $\Omega$; the eigenvectors of $\Omega$ determines the directions of the semi-axes [2]. Obviously, Bloch space is a special case of $\boldsymbol{A}=\vec{\pi}$. So, for arbitrary measurement vector $\boldsymbol{A}$, we have
$|\vec{\alpha}|^{2}\leq\frac{d-1}{d}\lambda_{\max}(\Omega)\;,$ (38)
where $\lambda_{\max}(X)$ is the maximal eigenvalue of matrix $X$. Finally, we prove
$\|\mathcal{C}\|_{\operatorname{tr}}\leq\kappa_{A}\kappa_{B}\;.$ (39)
Here, $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, similarly for $\kappa_{B}$.
The proof of Observation 1
If $\boldsymbol{A}$ has the same trace and the traceless part corresponds to a regular simplex, namely
$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (42)
where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e. $\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\},\mu,\nu=1,\cdots,d_{A}^{2}$, then we have $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=2\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$, whose matrix form is
$\Omega_{A}=\frac{2}{d_{A}^{2}-1}\left(d_{A}^{2}\mathds{1}-J_{d_{A}^{2}}\right)\;,$ (43)
where, $J_{d_{A}^{2}}$ is a $d_{A}^{2}$-dimensional all-ones matrix all whose elements are equal to one. On account of $J_{d_{A}^{2}}^{2}=d_{A}^{2}J_{d_{A}^{2}}$, it is readily to find eigenvalues of $J_{d_{A}^{2}}$, i.e. $d_{A}^{2}$ and 00 ($d_{A}^{2}-1$-fold degeneracy). Hence, we have $\lambda_{\max}(\Omega_{A})=\frac{2d_{A}^{2}}{d_{A}^{2}-1}$, and
$\kappa_{A}=\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\;,$ (44)
similarly for $\kappa_{B}$. So, for the separable state, we have
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;,$ (45)
Here, parameters $t_{A}$ and $t_{B}$ can be arbitrary real numbers. And next, we prove that, when $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}-1}}$ for $X=A,B$, Eq. 45 reproduces CCNR and ESIC (entanglement criterion via SIC-POVM) criteria, respectively. Firstly, we review the CCNR criterion. A bipartite state can be expressed as
$\rho=\frac{1}{4}\sum_{\mu,\nu}\chi_{\mu\nu}\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B}\;.$ (46)
Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$ and $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d_{X}^{2}-1}$ is generators of $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,B$; $\{\Pi_{\mu}^{X}\}_{\mu=0}^{d_{X}^{2}-1}$ satisfies the orthogonal relation $\mathrm{Tr}[\Pi_{\mu}^{X}\Pi_{\nu}^{X}]=2\delta_{\mu\nu}$. Making use of the Schmidt decomposition in operator space, we have
$\rho=\frac{1}{4}\sum_{\mu,\nu}\chi_{\mu\nu}\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B}=\sum_{k}\sigma_{k}\widetilde{\Pi}_{k}^{A}\otimes\widetilde{\Pi}_{k}^{B}\;.$ (47)
Here, $\sum_{k}\sigma_{k}=\frac{1}{2}\|\chi\|_{\mathrm{tr}}$ and $\{\widetilde{\Pi}_{k}^{A}\}$ and $\{\widetilde{\Pi}_{k}^{B}\}$ form orthonormal bases of the operator space. CCNR criterion says that, if $\rho$ is separable, then we have $\sum_{k}\sigma_{k}\leq 1$ [3, 4], that is, $\|\chi\|_{\mathrm{tr}}\leq 2$. If $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ for $X=A,B$, then we have
$M_{A}^{\mathrm{T}}M_{A}$ $=\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{1}\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{2}\\ \vdots&\vdots\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\frac{1}{\sqrt{d_{A}^{2}-1}}&\cdots&\frac{1}{\sqrt{d_{A}^{2}-1}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}\;.$ (54)
Here, $\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$ has been employed. Similarly, we have $M_{A}^{\mathrm{T}}M_{A}=\frac{d_{B}^{2}}{d_{B}^{2}-1}\mathds{1}$. So, if $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ for $X=A,B$, Eq. 45 gives $\|\mathcal{C}\|_{\mathrm{tr}}=\|M_{A}^{\mathrm{T}}\chi M_{B}\|_{\mathrm{tr}}=\frac{d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}\|\chi\|_{\mathrm{tr}}\leq\frac{2d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$, that is, $\|\chi\|_{\mathrm{tr}}\leq 2$.
And then, the ESIC criteria says that if $\rho$ is separable, then the correlation matrix of the normalized SIC-POVM $\{E_{\mu}^{A}\}$ and $\{E_{\mu}^{B}\}$ satisfies [5]
$\|\mathcal{P}\|_{\mathrm{tr}}\leq 1\;,$ (55)
where $\mathcal{P}_{\mu\nu}=\braket{E_{\mu}^{A}\otimes E_{\nu}^{B}}$, $E_{\mu}^{X}\geq 0$ and $\mathrm{Tr}[E_{\mu}^{X}E_{\nu}^{X}]=\frac{d_{X}\delta_{\mu\nu}+1}{2d_{X}},\mu,\nu=1,2,\cdots,d_{X}^{2}$ for $X=A,B$. It is noteworthy that the original assumption about the existence of SIC-POVM is not essential as shown in [6]. In fact, we can obtain ESIC criterion for arbitrary measurement vector $\boldsymbol{A}$ satisfying $\mathrm{Tr}[A_{\mu}A_{\nu}]=\alpha\frac{d\delta_{\mu\nu}+1}{2d},\mu,\nu=1,2,\cdots,d^{2}$ with nonzero real parameter $\alpha$. If $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}-1}}$ for $X=A,B$, we have $\mathrm{Tr}[A_{\mu}A_{\nu}]=\frac{t_{A}^{2}}{d}+2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{(2d_{A})^{2}}{d_{A}^{2}-1}\frac{d_{A}\delta_{\mu\nu}+1}{2d_{A}}$ and Eq. 45 gives
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\frac{4d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}\;,$ (56)
which reproduces the ESIC criterion due to $\braket{A_{\mu}\otimes B_{\nu}}=\frac{4d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}\braket{E_{\mu}^{A}\otimes E_{\nu}^{B}}$.
The vertex coordinates for regular simplex
There are many methods to generate the vertex coordinates for regular simplex, where we give a simple scheme for the reader’s convenience. Given the following $n$-dimensional unit vectors
$\vec{e}_{k}=(b_{1},b_{2},\cdots,b_{k-1},a_{k},0,\cdots,0)\;,$ (57)
and
$a_{k}=\frac{\sqrt{(n+1)(n-k+1)}}{\sqrt{n(n-k+2)}},b_{k}=-\frac{\sqrt{\ n+1}}{\sqrt{n(n-k+1)(n-k+2)}},k=1,2,\cdots,n+1,$ (58)
then $\{\vec{e}_{1},\vec{e}_{2}\cdots,\vec{e}_{n+1}\}$ constitutes a $n$-dimensional regular simplex. The following is an example for $n=8$,
$\left(\begin{array}[]{ccccccccc}1&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}\\ 0&\frac{3\sqrt{7}}{8}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}\\ 0&0&\frac{3}{2}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}\\ 0&0&0&\frac{\sqrt{15}}{4}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}\\ 0&0&0&0&\frac{3}{\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}\\ 0&0&0&0&0&\frac{3}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}\\ 0&0&0&0&0&0&\frac{\sqrt{3}}{2}&-\frac{\sqrt{3}}{4}&-\frac{\sqrt{3}}{4}\\ 0&0&0&0&0&0&0&\frac{3}{4}&-\frac{3}{4}\\ \end{array}\right)\;.$ (67)
Bound entangled states
Horodecki’s bound entangled states with white noise:
[…]
Here, $\rho_{\mathrm{H}}(t)$ is $3\times 3$ Horodecki’s bound entangled states [7], i.e.
$\rho_{\mathrm{H}}(s)=\frac{1}{8s+1}\left(\begin{matrix}s&0&0&0&s&0&0&0&s\\ 0&s&0&0&0&0&0&0&0\\ 0&0&s&0&0&0&0&0&0\\ 0&0&0&s&0&0&0&0&0\\ s&0&0&0&s&0&0&0&s\\ 0&0&0&0&0&s&0&0&0\\ 0&0&0&0&0&0&\frac{1}{2}(1+s)&0&\frac{1}{2}\sqrt{1-s^{2}}\\ 0&0&0&0&0&0&0&s&0\\ s&0&0&0&s&0&\frac{1}{2}\sqrt{1-s^{2}}&0&\frac{1}{2}(1+s)\end{matrix}\right)\;.$ (78)
Unextendible product bases (UPB) bound entangled states:
[…]
Here, $\rho_{\mathrm{UPB}}$ is bound entangled state constructed by UPB [8], i.e.
$\rho_{\mathrm{UPB}}=\frac{1}{4}\left(\mathds{1}-\sum_{i=0}^{4}\ket{\psi_{i}}\bra{\psi_{i}}\right)\;,$ (80)
and
[…]
Chessboard states [9]:
$\rho_{\mathrm{CB}}=\mathcal{N}\sum_{i=1}^{4}\ket{V_{i}}\bra{V_{i}}\;.$ | . (82)
Here, $\mathcal{N}=1/\sum_{i}\braket{V_{i}}{V_{i}}$ is normalization constant and the unnormalized vectors |$\ket{V_{i}}$ are
|$\ket{V_{1}}=\ket{m,0,ac/n;0,n,0;0,0,0}\;,$
|$\ket{V_{2}}=\ket{0,a,0;b,0,c;0,0,0}\;,$
|$\ket{V_{3}}=\ket{n,0,0;0,-m,0;ad/m,0,0}\;,$
|$\ket{V_{4}}=\ket{0,b,0;-a,0,0;0,d,0}\;.$ (83)
Chessboard states contain six real parameters $m,n,a,b,c,d$. In the main text, we test Observation 1 by randomly generated $50000$ chessboard states, where the six parameters are drawn independently from the standard normal distribution.
The proof of Theorem 2
By virtue of Bloch representation, the correlation matrix can be written as $\mathcal{C}=M_{A}^{\mathrm{T}}\chi M_{B}$, so we have
$\|\mathcal{C}\|_{\mathrm{tr}}=\mathrm{Tr}\left[\sqrt{\mathcal{C}^{\mathrm{T}}\mathcal{C}}\right]=\mathrm{Tr}\left[\sqrt{M_{B}M_{B}^{\mathrm{T}}\chi^{\mathrm{T}}M_{A}M_{A}^{\mathrm{T}}\chi}\right]\;.$ (84)
Given arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, we have
$M_{A}=\left(\begin{matrix}\frac{t_{1}^{A}}{\sqrt{2d_{A}}}&\frac{t_{2}^{A}}{\sqrt{2d_{A}}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;,\;M_{B}=\left(\begin{matrix}\frac{t_{1}^{B}}{\sqrt{2d_{B}}}&\frac{t_{2}^{B}}{\sqrt{2d_{B}}}&\cdots\\ \vec{b}_{1}&\vec{b}_{2}&\cdots\end{matrix}\right)\;,$ (89)
and
$M_{A}M_{A}^{\mathrm{T}}=\left(\begin{matrix}\frac{(\vec{t}^{A})^{2}}{2d_{A}}&0\\ 0&\mathcal{A}\end{matrix}\right)=\frac{(\vec{t}^{A})^{2}}{2d_{A}}\oplus\mathcal{A}\;,$ (92)
where $\mathcal{A}=\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}$. Similarly, we have $M_{B}M_{B}^{\mathrm{T}}=\frac{(\vec{t}^{B})^{2}}{2d_{B}}\oplus\mathcal{B}$ and $\mathcal{B}=\sum_{\mu}\vec{b}_{\mu}\vec{b}_{\mu}^{\mathrm{T}}$. For the normal form states, we have $\chi=\frac{2}{\sqrt{d_{A}d_{B}}}\oplus\widetilde{\mathcal{T}}$. And then,
$\|\mathcal{C}\|_{\mathrm{tr}}$ $=\mathrm{Tr}\left[\sqrt{M_{B}M_{B}^{\mathrm{T}}\chi^{\mathrm{T}}M_{A}M_{A}^{\mathrm{T}}\chi}\right]$ (93)
$=\mathrm{Tr}\left[\sqrt{\frac{(\vec{t}^{A})^{2}(\vec{t}^{B})^{2}}{d_{A}^{2}d_{B}^{2}}\oplus\mathcal{B}\widetilde{\mathcal{T}}^{\mathrm{T}}\mathcal{A}\widetilde{\mathcal{T}}}\right]$ (94)
$=\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}+\mathrm{Tr}\left[\sqrt{\mathcal{A}\widetilde{\mathcal{T}}\mathcal{B}\widetilde{\mathcal{T}}^{\mathrm{T}}}\right]\;.$ (95)
So, if a quantum state $\rho$ is separable, then we have in its normal form $\tilde{\rho}$
$\mathrm{Tr}\left[\sqrt{\mathcal{A}\widetilde{\mathcal{T}}\mathcal{B}\widetilde{\mathcal{T}}^{\mathrm{T}}}\right]\leq\kappa_{A}\kappa_{B}-\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}\;.$ (96)

### v2

@@
 Abstract
-We propose in this work a practical approach, by virtue of correlation matrices of the generic observables, to study the long lasting tough issue of quantum separability.
-Some general separability conditions are set up through constructing a measurement-induced Bloch space.
-In essence, these conditions are established due to the self constraint in the space of quantum states.
-The new approach can not only reproduce many of the prevailing entanglement criteria, but also lead to even stronger results and manifest superiority for some bound entangled states.
-Moreover, as a by product, the new criteria are found directly transformable to the entanglement witness operators.
-Introduction
-Quantum entanglement captures the most enigmatic and deepest insights into quantum theory, and has deep connections with other branches of physics [1].
-In quantum information theory, how to distinguish an entangled state from a separable state is a fundamental and significant problem, say the so-called separability problem [2, 3].
-It is well-known that the separability problem tends to be a NP-hard issue with the increase of dimension, even for the bipartite system [4], to develop a practical and experimental friendly separability criterion is pivotal in the establishment of a complete entanglement theory [2, 3].
-In past decades, numerous methods have been put forward to attack it, including positive map theory and entanglement witness [5, 6, 7], the computable cross norm or realignment (CCNR) criterion [8, 4, 3], correlation matrix method [11, 12, 13, 5], local uncertainty relation (LUR) criteria [15, 16, 17, 18, 19, 20], covariance matrix [21, 22], quantum Fisher information (QFI) method [23, 24, 25], moment method [26, 27, 28, 29], etc.
-In this Letter, we develop a framework to derive separability criterion for finite-dimensional systems in terms of the correlation matrix.
-To realize it, a measurement-induced, generalized Bloch space for the arbitrary observables is introduced, and by which a generic separability criterion is formulated.
-The key point here is that the correlation matrix of the separable state is limited by the size of the generalized Bloch space.
-The powerfulness of this approach can be realized by noticing that in which the de Vicente’s correlation matrix criterion [11], CCNR criterion [4, 3], the correlation matrix criterion based on symmetric informationally complete positive operator-valued measures (SIC-POVM) [5] etc.
-can all be reproduced transparently, and even stronger criterion may be obtained.
-Moreover, the new scenario also provides a concise method to construct an entanglement witness operator via generic observables.
-Examples are provided to show the availability of bound entangled state detection.
-Bloch representation of observables and density matrix
-The notations to be employed in the discussion go as follows.
-The observable is denoted by a capital letter, such as $A$, and a boldface capital letter $\boldsymbol{A}$ implies a measurement vector containing certain observables, i.e.
-$\boldsymbol{A}=(A_{1},A_{2},\cdots)$.
-The density matrix $\rho$ of a $d$-level quantum system is formulated in Bloch representation, to wit [30, 31, 32, 33, 35, 34]
+We propose in this work a practical approach to address the longstanding and challenging problem of quantum separability, leveraging the correlation matrices of generic observables.
+General separability conditions are obtained by dint of constructing the measurement-induced Bloch space, which in essence come from the intrinsic constraints in the space of quantum state.
+The novel approach can not only reproduce various established entanglement criteria, it may as well brings about some new results, possessing obvious advantages for certain bound entangled states and the high dimensional Werner states.
+Moreover, it is found that criteria obtained in our approach can be directly transformed into entanglement witness operators.
+1 Introduction
+Quantum entanglement captures the most enigmatic and deepest insights of quantum theory, and has a broad connections with other branches of physics [1].
+In quantum information theory, distinguishing entangled states from separable states appears to be a fundamental and challenging issue, commonly referred to as the separability problem [2, 3].
+It is widely recognized that the separability problem becomes NP-hard as the dimension increases, even in bipartite systems [4].
+Therefore, developing a practical and experimentally feasible separability criterion is crucial for establishing a comprehensive entanglement theory [2, 3].
+In past decades, a multitude of approaches have been proposed to address this issue, including positive map theory and entanglement witnesses [5, 6, 7], spin squeezing inequalities [8, 9, 10, 11], the computable cross norm or realignment (CCNR) criterion [12, 13, 14], correlation matrix method [15, 16, 17, 18, 19], local uncertainty relation (LUR) criteria [20, 21, 22, 23, 24, 25], covariance matrix [26, 27], quantum Fisher information (QFI) method [28, 29, 30, 31], moment method [32, 33, 34, 35], etc.
+In this paper, we propose a comprehensive framework to establish a separability criterion for finite-dimensional quantum systems using the correlation matrix.
+To achieve this, we introduce a measurement-induced generalized Bloch space for arbitrary observables, which enables us to formulate a generic separability criterion.
+A critical observation is that the correlation matrix of separable states is constrained by the scale parameter of the generalized Bloch space.
+The versatility and strength of this approach are evident from its ability to reproduce several established criteria, such as de Vicente’s correlation matrix criterion [15], the CCNR criterion [13, 14], the correlation matrix criterion based on symmetric informationally complete positive operator-valued measures (SIC-POVM) [18] and Sarbicki et al.’s criterion [19], while the potential to yield some other even stronger results.
+Additionally, this framework provides a streamlined method for constructing entanglement witness operators using generic observables.
+We provide examples demonstrating the effectiveness of detecting bound entangled states and the high-dimensional Werner states.
+2 Bloch representation of observables and density matrix
+Before proceeding, we introduce the necessary notations.
+An observable is denoted by a capital letter, such as $A$, while a boldface capital letter $\boldsymbol{A}$ represents an $m$-tuple of observables, i.e., $\boldsymbol{A}=(A_{1},A_{2},\cdots)$.
+The density matrix $\rho$ of a $d$-level quantum system is expressed in the Bloch representation, specifically [36, 37, 38, 39, 40, 41]
 $\rho=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{r}\cdot\vec{\pi}\;.$ (1)
@@
 where $t_{\mu}=\mathrm{Tr}[A_{\mu}]$ and $a_{\mu\nu}=(\vec{a}_{\mu})_{\nu}=\mathrm{Tr}[A_{\mu}\pi_{\nu}]/2$.
-And then, a measurement vector $\boldsymbol{A}$ can be parameterized as a real matrix $M$ in Bloch representation
-$\boldsymbol{A}\rightarrow M_{A}\equiv\left(\begin{matrix}\frac{t_{1}}{\sqrt{2d}}&\frac{t_{2}}{\sqrt{2d}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;.$ (5)
+And then, an $m$-tuple $\boldsymbol{A}$ of observables can be parameterized as a real matrix $M$ in Bloch representation
+$\boldsymbol{A}\rightarrow M_{A}\equiv\left[\begin{matrix}\frac{t_{1}}{\sqrt{2d}}&\frac{t_{2}}{\sqrt{2d}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right]\;.$ (3)
 Note, here $\vec{a}_{\mu}$ are column vectors.
-Correlation matrix of separable state and measurement-induced Bloch space
-Quantum entanglement exhibits inherent non-classical correlation in many-body systems, even far apart from each other, which was initially recognized by Einstein, Podolsky and Rosen [36] and Schrödinger [37].
-Based on correlation function [38] Bell inequality provides a practical way for the first time to test the entanglement.
-In the family of Bell inequality, the Clauser-Horne-Shimony-Holt (CHSH) inequality [39]
-$\left|\braket{A\otimes B}-\braket{A\otimes B^{\prime}}+\braket{A^{\prime}\otimes B}+\braket{A^{\prime}\otimes B^{\prime}}\right|\leq 2\;$ (6)
-might be the most prominent one, which tends to be experiment-operational friendly.
+3 Correlation matrix of separable state and measurement-induced Bloch space
+Quantum entanglement exhibits inherent non-classical correlations in many-body systems, even when the particles are far apart.
+This phenomenon was initially recognized by Einstein, Podolsky, and Rosen [42], as well as by Schrödinger [43].
+Based on correlation functions, Bell inequalities provide a practical method to test entanglement for the first time [44].
+Among the family of Bell inequalities, the Clauser-Horne-Shimony-Holt (CHSH) inequality [45]
+$\left|\!\braket{A\otimes B}\!\!-\!\!\braket{A\otimes B^{\prime}}\!\!+\!\!\braket{A^{\prime}\otimes B}\!\!+\!\!\braket{A^{\prime}\otimes B^{\prime}}\!\right|\!\leq\!2$ (4)
+is perhaps the most prominent.
 Here, $A(A^{\prime})$ and $B(B^{\prime})$ are pairs of dichotomic observables with eigenvalues of ±$\pm 1$ measured by the parties $A$ and $B$, respectively.
 The violation of CHSH inequality and the predictions of quantum theory were first confirmed experimentally by [46, 47, 48], then later more convincingly by some loophole-free tests [49, 50, 51, 52].
-The violation of CHSH inequality implies the entanglement and exclusion of local hidden variable scenario for the description of the correlation system.
-However, in reverse, Bell inequalities fail to recognize all the entangled states [47].
-Hence, to explore and exploit the nature of entanglement, it is essential to find even stronger criteria for the determination of separable state.
-To this end, we define the following correlation matrix entry for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$
+Although the violation of the CHSH inequality certifies entanglement and excludes the possibility of a local hidden variable description for correlated systems, Bell inequalities inherently fail to detect all entangled states [53].
+Therefore, to fully explore and utilize the nature of entanglement, it is essential to establish even stronger criteria for determining separable states.
+To this end, we define the following correlation matrix entries for arbitrary $m$-tuples $\boldsymbol{A}$ and $\boldsymbol{B}$
 $\mathcal{C}_{\mu\nu}\equiv\braket{A_{\mu}\otimes B_{\nu}}\;.$ (5)
-Generally speaking, a state is separable iff it is a convex combination of product states, i.e.
-$\rho_{\mathrm{sep}}=\sum_{i}q_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$, where $\{\rho_{i}^{A}\}$ and $\{\rho_{i}^{B}\}$ are two pure subsystems.
-And, hence the corresponding correlation matrix for a separable state can write
+A state is separable iff it is a convex combination of product states, i.e.
+$\rho_{\mathrm{sep}}=\sum_{i}q_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$, where $\{\rho_{i}^{A}\}$ and $\{\rho_{i}^{B}\}$ are two pure subsystems [53].
+Hence the correlation matrix of a separable state can be written as
 $\mathcal{C}=\sum_{i}q_{i}\left(\begin{matrix}\braket{A_{1}}_{i}\\ \braket{A_{2}}_{i}\\ \vdots\end{matrix}\right)\left(\begin{matrix}\braket{B_{1}}_{i},\braket{B_{2}}_{i},\cdots\end{matrix}\right)\;,$ (6)
@@
 Eq.
-12 inspires us to define a measurement-induced Bloch vector(MIBV) $\vec{\alpha}$ for measurements $\boldsymbol{A}$, of which
+6 inspires us to define a measurement-induced Bloch vector (MIBV) $\vec{\alpha}$ for $\boldsymbol{A}$, given by
 $\alpha_{\mu}\equiv\mathrm{Tr}[\rho\bar{A}_{\mu}]=\vec{r}\cdot\vec{a}_{\mu}\;$ (7)
-with $\bar{A}_{\mu}=\vec{a}_{\mu}\cdot\vec{\pi}$ denoting the traceless part of observable $A_{\mu}$.
-And, various MIBV $\vec{\alpha}$ span the corresponding measurement-induced Bloch space, signified as $\mathcal{B}(\boldsymbol{A})$, which has been applied to the study of certainty and uncertainty relations [1].
-The correlation matrix of a separable state can be reformulated by MIBVs, like
+where $\bar{A}_{\mu}=\vec{a}_{\mu}\cdot\vec{\pi}$ denotes the traceless part of the observable $A_{\mu}$.
+Various MIBV $\vec{\alpha}$ span the corresponding measurement-induced Bloch space (MIBS), denoted as $\mathcal{B}(\boldsymbol{A})$, which is defined as
+$\mathcal{B}(\!\boldsymbol{A}\!)\!=\!\{\vec{\alpha}\!=\!(\braket{\!\bar{A}_{1}\!}_{\rho},\!\cdots,\!\braket{\!\bar{A}_{m}\!}_{\rho})^{\mathrm{T}}\!|\rho\in\!\mathcal{S}(\mathcal{H})\!\}\;.$ (8)
+Here, $\mathcal{S}(\mathcal{H})$ is the set of density matrices.
+It is noteworthy that $\mathcal{B}(\boldsymbol{A})$ is also referred to as the joint numerical range (JNR) of observables $\bar{A}_{1},\cdots,\bar{A}_{m}$ [54, 55], which has been extensively studied in the context of uncertainty relations, quantum entanglement, and more [25, 56].
+In particular, the quantity
+$\beta\equiv\sup_{\rho}\sum_{k=1}^{m}\braket{\bar{A}_{k}}_{\rho}^{2}=\max_{\vec{x}\in\mathcal{B}(\boldsymbol{A})}|\vec{x}|^{2}$ (9)
+is a critical parameter of JNR or MIBS, closely related to state-independent uncertainty relation [57, 58].
+In the following sections, we will establish an explicit relationship between $\beta$ and the quantum separability problem.
+The correlation matrix of a separable state can be reformulated using MIBVs, as follows:
 $\mathcal{C}=\sum_{i}q_{i}(\vec{t}^{A}/d_{A}+\vec{\alpha}_{i})(\vec{t}^{B}/d_{B}+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (10)
-Here, $\vec{t}^{A}$ and $\vec{t}^{B}$ are column vectors consisting of traces of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, respectively; $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are MIBVs of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$; $\mathrm{T}$ denotes the transpose of a given matrix.
-The separability criteria
-A family of sufficient conditions for entanglement of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$ goes as follows, with proof given in the Supplement:
+Here, $\vec{t}^{A}$ and $\vec{t}^{B}$ are column vectors consisting of the traces of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$, respectively; $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are MIBVs of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$; and $\mathrm{T}$ denotes the transpose of a given matrix.
+4 The separability criteria
+With this, we are now ready to present the main results of this work, with proof given in the Appendix A:
 Theorem 1.
-Let $\mathcal{C}$ be a correlation matrix for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, if the corresponding state $\rho$ is separable, then
+Let $\mathcal{C}$ be a correlation matrix for arbitrary $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, if the corresponding state $\rho$ is separable, then
 $\|\mathcal{C}\|_{\operatorname{tr}}\leq\kappa_{A}\kappa_{B}\;.$ (11)
 Otherwise, $\rho$ will be an entangled state.
-Here, $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$ is the trace norm of matrix $X$; $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ with $\lambda_{\max}(X)$ being the maximal eigenvalue of matrix $X$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, and similarly for $\kappa_{B}$.
-Next, we illustrate how Theorem 1 works. For a bipartite system, we have the following Bloch representation
+Here, $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$ is the trace norm of matrix $X$; $\kappa_{A}=\sqrt{(\vec{t}^{\,A})^{2}/d_{A}^{2}+\beta_{A}}$ with $\beta_{A}=\sup_{\rho}\sum_{k=1}^{m}\braket{A_{k}}_{\rho}^{2}$, and similarly for $\kappa_{B}$.
+Note, Theorem 1 provides an alternative approach to determining the separability of quantum states, different from the covariance matrix (CM) criteria [26, 27].
+The CM criteria make use of the concavity and the positive definitiveness of CM, while Theorem 1 stems from the geometry property of the MIBS, highlighting the independence of the covariance matrix and correlation matrix methods.
+Next, we illustrate how Theorem 1 works.
+For a bipartite system, we have the following Bloch representation
 $\rho$ $=\frac{1}{4}\sum_{\mu,\nu}\chi_{\mu\nu}\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B}$ (12)
 $=\frac{1}{d_{A}d_{B}}\mathds{1}\otimes\mathds{1}+\frac{1}{2d_{B}}\vec{a}\cdot\vec{\pi}^{A}\otimes\mathds{1}+\frac{1}{2d_{A}}\mathds{1}\otimes\vec{b}\cdot\vec{\pi}^{B}+\frac{1}{4}\sum_{\mu,\nu}\mathcal{T}_{\mu\nu}\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}\,.$ (13)
-Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$; $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d^{2}-1}$ are generators of $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,\ B$; the coefficient matrix $\chi_{\mu\nu}=\mathrm{Tr}[\rho(\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B})]$; $a_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\mathds{1}]$; $b_{\mu}=\mathrm{Tr}[\rho\mathds{1}\otimes\pi_{\mu}^{B}]$; $\mathcal{T}_{\mu\nu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}]$; and $\chi_{00}=\frac{2}{\sqrt{d_{A}d_{B}}}$ since $\mathrm{Tr}[\rho]=1$.
-The correlation matrix can then be reexpressed as
+Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$; $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d^{2}-1}$ are the generators of the $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,B$; coefficient matrix is given by $\chi_{\mu\nu}=\mathrm{Tr}[\rho(\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B})]$; $a_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\mathds{1}]$; $b_{\mu}=\mathrm{Tr}[\rho\mathds{1}\otimes\pi_{\mu}^{B}]$; $\mathcal{T}_{\mu\nu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}]$; and $\chi_{00}=\frac{2}{\sqrt{d_{A}d_{B}}}$ since $\mathrm{Tr}[\rho]=1$.
+The correlation matrix can then be expressed as
 $\mathcal{C}=M_{A}^{\mathrm{T}}\chi M_{B}\;$ (14)
-with $M_{A}$ and $M_{B}$ being the Bloch representations of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ defined in Eq.
-5.
-Equipped with this, one may notice that the traceless part of the measurement $\boldsymbol{A}$ can be viewed as the vertex of a polygon or polyhedron in $(d^{2}-1)$-dimensional space, which is enlightening for further processing of Theorem 1.
-Consider the simplest situation $t_{\mu}^{A}=0$, and $\vec{a}_{1}=\vec{0}$, $\vec{a}_{\mu}=(0,\cdots,0,1,0,\cdots,0)^{\mathrm{T}}$, $\mu=2,\cdots,d_{A}^{2}$, where all components of $\vec{a}_{\mu}$ equal to 00, except the $\mu$th equals to $1$.
-It is easy to find $\lambda_{\max}(\Omega_{A})=2$ (similarly for $B$) and the separability condition
+where $M_{A}$ and $M_{B}$ are the Bloch representations of $\boldsymbol{A}$ and $\boldsymbol{B}$ defined in Eq.
+3.
+With this framework, one may observe that the traceless part of the measurement $\boldsymbol{A}$ can be interpreted as the vertices of a polygon or polyhedron in $(d^{2}-1)$-dimensional space, which provides valuable insights for the further analysis of Theorem 1.
+Let us consider the simplest scenario where $t_{\mu}^{A}=0$, $\vec{a}_{1}=\vec{0}$, and $\vec{a}_{\mu}=(0,\cdots,0,1,0,\cdots,0)^{\mathrm{T}}$ for $\mu=2,\cdots,d_{A}^{2}$.
+Here, all components of $\vec{a}_{\mu}$ are zero except for the $(\mu-1)$th component which equals 1.
+It is straightforward to determine that $\beta_{A}=\frac{2(d_{A}-1)}{d_{A}}$ (analogously for $\beta_{B}$).
+The separability condition can be expressed as:
 $\|\mathcal{C}\|_{\mathrm{tr}}=\|\mathcal{T}\|_{\mathrm{tr}}\leq\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}\;.$ (15)
-Here, $\mathcal{T}=\mathrm{Tr}[\rho(\pi_{\mu}\otimes\pi_{\nu})]$. Note, Eq.(19) simply reproduces the de Vicente’s correlation matrix criterion [11].
-Moreover, if we let $t_{1}^{A}=\sqrt{2d}h_{A},t_{1}^{B}=\sqrt{2d}h_{A},t_{\mu}^{A}=t_{\mu}^{B}=0,\mu\neq 1$, and take $\{\vec{a}_{\mu}\}$ and $\{\vec{b}_{\mu}\}$ the same as in above, we obtain
-$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{\frac{2(d_{A}-1+h_{A}^{2})}{d_{B}}}\sqrt{\frac{2(d_{B}-1+h_{B}^{2})}{d_{B}}}\;,$ (20)
-which is exactly the Sarbicki et al.’s criterion [6]. Whereas, if $\boldsymbol{A}$, and similarly for $\boldsymbol{B}$, has the same trace and its traceless part corresponds to a regular simplex, viz.
-$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (23)
-where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e.
-$\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\}$ for $\mu,\nu=1,\cdots,d_{A}^{2}$, a family of separability criteria can be readily obtained (the proof of separability criteria and construction of regular simplex vertex coordinates are given in Supplements):
+Notably, Eq.(15) simply reproduces the de Vicente’s correlation matrix criterion [15].
+Furthermore, by setting $t_{1}^{A}=\sqrt{2d_{A}}h_{A}$, $t_{1}^{B}=\sqrt{2d_{B}}h_{B}$, and $t_{\mu}^{A}=t_{\mu}^{B}=0$ for $\mu\neq 1$, and choosing $\{\vec{a}_{\mu}\}$ and $\{\vec{b}_{\mu}\}$ as previously defined, we obtain
+$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{\frac{2(d_{A}-1+h_{A}^{2})}{d_{A}}}\sqrt{\frac{2(d_{B}-1+h_{B}^{2})}{d_{B}}},$
+which is precisely the Sarbicki et al.’s criterion [19].
+Additionally, if $\boldsymbol{A}$ (and similarly for $\boldsymbol{B}$) has the same trace and its traceless part corresponds to a regular simplex, specifically,
+$M_{A}=\left[\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\;,$ (16)
+where $\{\vec{a}_{\mu}\}$ forms a regular $(d_{A}^{2}-1)$-simplex, i.e., $\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$ for $\mu,\nu=1,\ldots,d_{A}^{2}$, a family of separability criteria can be readily derived.
+The proof and the construction of the regular simplex vertex coordinates are provided in Appendices B and C.
 Observation 1.
-Given measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$ and traceless parts correspond respectively to a regular simplex.
-The correlation matrix of separable state $\rho$ then satisfy
+Given $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$ and traceless parts correspond respectively to a regular simplex.
+The correlation matrix of separable state $\rho$ then satisfies
 $\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;,$ (17)
 otherwise, $\rho$ is entangled. Here, parameters $t_{A}$ and $t_{B}$ are arbitrary real numbers.
-The 1 represents a family of separability conditions for different $t_{A}$ and $t_{B}$, many of existing criteria appear to be its particular cases.
-When $t_{X}$ takes $\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $\sqrt{\frac{2d_{X}}{d_{X}-1}}$ ($X=A,B$), it gives out the well-known CCNR [50, 4] and entanglement criterion via SIC-POVM (ESIC) criteria [5], respectively.
-To exhibit more, following we compare 1 with some symbolic criteria under certain quantum states, the Horodecki’s bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$, unextendible product bases (UPB) bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and chessboard states $\rho_{\mathrm{CB}}$.
-The exact forms of these quantum states are shown in the Supplemental Material.
-In Fig.
-1 we show and compare 1, CCNR and ESIC criteria under three typical classes of bound entangled states, where larger $t$ witnesses higher robustness against white noise and hence at which larger portion of entangled state may be detected.
-Fig.
-1(a) exhibits Horodecki’s bound entangled states with white noise and Fig.
-1(b) explains the detection results for UPB state with white noise (peru solid line) and $50000$ randomly generated chessboard state (blue dashed line).
-Note, as shown in the figure, 1 reproduces the CCNR and ESIC criteria as two special cases when $t=\sqrt{3}/2$ and $t=\sqrt{3}$, respectively.
+The 1 represents a family of separability conditions for various values of $t_{A}$ and $t_{B}$, with many existing criteria appearing as particular cases of this family.
+Specifically, when $t_{X}$ takes the values $\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $\sqrt{\frac{2d_{X}}{d_{X}-1}}$ ($X=A,B$), it yields the well-known CCNR criterion [59, 13] and the entanglement criterion via SIC-POVM (ESIC) [18], respectively.
+For clarity and convenience, we have summarized these results in Table 1.
+Table 1: Here, we summarize the relations between our criteria and these mentioned in the main text without local filtering transformation, in which CCNR and ESIC are derived from 1 with $t=\sqrt{\frac{2d}{d^{2}-1}}$ and $t=\sqrt{\frac{2d}{d-1}}$, respectively.
+Criteria $M_{A}/M_{B}$ $\kappa_{A}\kappa_{B}$
+Vicente [15] $\left[\begin{smallmatrix}0&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}$
+Sarbicki [19] $\left[\begin{smallmatrix}h&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\sqrt{\frac{4(d_{A}-1+h_{A}^{2})(d_{B}-1+h_{B}^{2})}{d_{A}d_{B}}}$
+This work $\left[\begin{smallmatrix}\frac{t}{\sqrt{2d}}&\cdots&\frac{t}{\sqrt{2d}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}$
+CCNR [59, 13] $\left[\begin{smallmatrix}\frac{1}{\sqrt{d^{2}-1}}&\cdots&\frac{1}{\sqrt{d^{2}-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{2d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
+ESIC [18] $\left[\begin{smallmatrix}\frac{1}{\sqrt{d-1}}&\cdots&\frac{1}{\sqrt{d-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{4d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
+To elaborate further, we compare 1 with several symbolic criteria under specific quantum states: the Horodecki’s bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$, unextendible product bases (UPB) bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$, and chessboard states $\rho_{\mathrm{CB}}$.
+The exact forms of these quantum states are shown in the Appendix E.
 (a) Horodecki’s bound entangled states with white noise
 (b) UPB and chessboard states
-Figure 1: Entanglement detection criteria.
+Figure 1: Entanglement detection for three classes of bound entangled states.
 (a) Horodecki’s $3\times 3$ bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$.
 (b) UPB bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and $50000$ randomly generated chessboard states $\rho_{\mathrm{CB}}$.
-Here, we simply take $t_{A}=t_{B}=t$.
 The larger $t$ detects the more entangled states as shown in (a) and (b), in which $t=\sqrt{3}/2$ and $t=\sqrt{3}$ correspond to CCNR and ESIC criteria, respectively.
 Diagrams indicate explicitly that the criterion given in Observation 1 outperform CCNR and ESIC criteria under the three typical classes of bound entangled states.
-Encouraged by the results in Fig.1, one may naturally ask whether larger $t$ means a more powerful criterion, or, specifically, whether the ESIC criterion is stronger than the CCNR or not, which was conjectured to be true in Ref.
+In Fig.
+1, we compare the performance of 1, the CCNR criterion, and the ESIC criterion for three classes of bound entangled states.
+Here, a larger value of $t$ indicates greater robustness against white noise, enabling the detection of a larger portion of entangled states.
+Fig.
+1(a) presents results for Horodecki’s bound entangled states with white noise, while Fig.
+1(b) details the detection results for the UPB state with white noise (solid peru line) and $50000$ randomly generated chessboard states (blue dashed line).
+Notably, as illustrated in the figure, 1 reproduces the CCNR and ESIC criteria as two special cases when $t=\sqrt{3}/2$ and $t=\sqrt{3}$, respectively.
+When the two involved subsystems share the same dimension, an enhanced separability condition is readily established (see Appendix D for the proof):
+Observation 2.
+Given $\boldsymbol{A}=\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{\mu}^{B}=t$ and traceless parts correspond to a regular simplex.
+The correlation matrix of the $d\times d$ separable state $\rho$ then satisfies
+$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;,$ (18)
+and
+$\mathrm{Tr}[\mathcal{C}]\geq t^{2}-\frac{2d}{d^{2}-1}\;,$ (19)
+otherwise, $\rho$ is entangled. Here, $t$ is arbitrary real parameter.
+We now compare the criterion proposed by Sarbicki et al.
+[19] with 2.
+To this end, we first consider the random $2\times 2$, $3\times 3$ and $5\times 5$ states with the Hilbert-Schmidt distribution [^1] If a square random matrix $A$ whose entries are drawn independently according to the standard complex Gaussian distribution, then the density matrix $\rho=AA^{\dagger}/\mathrm{Tr}[AA^{\dagger}]$ obeys the Hilbert-Schmidt distribution [18]..
+In Fig.
+2, we plot the fractions of entangled states identified by Eq.
+18 (peru solid line) and the criterion proposed by Sarbicki et al.
+(blue solid line) for 50000 randomly generated $2\times 2$, $3\times 3$ and $5\times 5$ states following the Hilbert-Schmidt distribution.
+As the parameters $t$ and $h$ increase, the two inequalities tend to align, detecting the same fractions of entangled states.
+Therefore, we conjecture that the criterion proposed by Sarbicki et al.
+and 1 are equivalent when the entire parameter spaces are explored.
+Notably, a significantly larger percentage of entangled states is detected in the $3\times 3$ case compared to other dimensions, as highlighted in [18], underscoring the uniqueness of dimension $3$.
+Figure 2: Entanglement detection for the randomly generated $2\times 2$, $3\times 3$ and $5\times 5$ states with the Hilbert-Schmidt distribution.
+With the parameters $t$ or $h$ increasing, the two inequalities tend to be agreeable and detect the same fractions.
+Next, we specifically consider arbitrary dimensional Werner states [53, 60]
+$\rho_{\mathrm{W}}\!=\!\frac{1}{d^{2}}\mathds{1}\otimes\mathds{1}\!+\!\frac{1}{4}\sum_{\mu=1}^{d^{2}-1}\frac{2(d\phi-1)}{d(d^{2}-1)}\pi_{\mu}\otimes\pi_{\mu}\;,$ (20)
+where the parameter range is $\phi\in[-1,1]$.
+The Werner states are separable if and only if $\phi\geq 0$ [53].
+For the arbitrary dimensional Werner states, Sarbicki et al.’s criteria give $|d\phi-1|\leq\sqrt{d-1+h_{A}^{2}}\sqrt{d-1+h_{B}^{2}}$, if $\rho_{\mathrm{W}}$ is separable.
+The optimal inequality constraint is achieved by $\min_{\{h_{A},h_{B}\}}\sqrt{d-1+h_{A}^{2}}\sqrt{d-1+h_{B}^{2}}=d-1$, which recognizes the entangled range $\phi\in[-1,-\frac{d-2}{d})$.
+Notably, the intervals of entanglement detectable by Sarbicki et al.’s criteria tend to diminish to zero as the dimensionality increases.
+A straightforward calculation demonstrates that Eq.
+19 serves as a necessary and sufficient condition applicable across the entire parameter space $\phi\in[-1,0)$ for the arbitrary dimensional Werner states.
+5 Enhanced separability criteria and local filtering transformation
+Encouraged by the results in Fig.1, one might naturally ask whether a larger $t$ indicates a more powerful criterion.
+Specifically, the question arises as to whether the ESIC criterion is stronger than the CCNR criterion, a conjecture proposed in Ref.
 [18].
 Unfortunately, Eq.
-45 can not tell, since the left hand side (lhs) of it contains as well parameters $t_{A}$ and $t_{B}$.
-Nevertheless, we find this question can be readily addressed under the local filtering transformation (LFT), which preserves the separability of a given state, i.e.
+17 cannot tell more, since the left hand side (LHS) of it contains as well parameters $t_{A}$ and $t_{B}$.
+Nevertheless, it is found that this question can be readily addressed under the local filtering transformation (LFT), which preserves the separability of a given state, i.e.
 $\rho\xmapsto{}\tilde{\rho}=(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}/\mathrm{Tr}[(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}]$ with $F_{A/B}$ the arbitrary invertible matrices [61].
@@
 $\tilde{\rho}=\frac{1}{d_{A}d_{B}}\mathds{1}\otimes\mathds{1}+\frac{1}{4}\sum_{\mu,\nu}\widetilde{\mathcal{T}}_{\mu\nu}\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}\;.$ (21)
-By dint of quantum states formulated in form of Eq. 25 the bipartite separability problem [52] will be solved. Concretely, we have the following theorem (see Supplemental Material for proof):
+It is proved that all $d_{A}\times d_{B}$ states with local ranks $n\leq d_{A}$ and $m\leq d_{B}$ are either reducible to $n\times m$ bipartite states with full local ranks, or entangled in Ref.
+[62].
+Therefore we only need to focus on the separability problem of quantum states with full local ranks, which has been depicted in the flow diagram of Fig.
+3.
+Figure 3: The flow diagram of entanglement detection. The bipartite separability problem reduces to the bipartite states with full local ranks and further normal form states.
+In this sense, the bipartite separability problem can be addressed through quantum states expressed in the form of Eq.
+21.
+Specifically, we present the following theorem (see Appendix F for the proof):
 Theorem 2.
@@
 If $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, then we readily have $\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}$ and the following Observation:
-Observation 2.
+Observation 3.
 Given measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$, and traceless parts correspond respectively to a regular simplex.
@@
 Now that the LHS of Eq.
-27 is independent of parameters $t_{A},t_{B}$, one can then strictly compare it with other criteria for different $t_{A}$ and $t_{B}$.
+23 is independent of parameters $t_{A},t_{B}$, we can strictly compare it with other criteria for different values $t_{A}$ and $t_{B}$.
 Notice that the minimal value of $\kappa(t_{A},t_{B})$, i.e.
@@
 [26] when $d_{A}=d_{B}$.
-Furthermore, the CCNR and ESIC criteria will reexhibit respectively while we take $\kappa(t_{A},t_{B})=2-2/\sqrt{d_{A}d_{B}}$ and $4-2\sqrt{(d_{A}+1)(d_{B}+1)/d_{A}d_{B}}$, which also implies that in normal form the ESIC criterion is stronger than CCNR and they coincide as $d_{A}=d_{B}$.
-Relations to entanglement witness
+After LFT, the CCNR and ESIC criteria correspond to $\kappa(t_{A},t_{B})=2-2/\sqrt{d_{A}d_{B}}$ and $4-2\sqrt{(d_{A}+1)(d_{B}+1)/d_{A}d_{B}}$ respectively.
+This indicates that in normal form, the ESIC criterion is stronger than CCNR, with both coinciding when $d_{A}=d_{B}$.
+In Table 2, we present a summary of how to obtain de Vicente’s criterion, Sarbicki et al.’s criterion, CCNR, ESIC, and Proposition 6 in [26] after applying local filtering transformations, highlighting the explicit relationships among them.
+Table 2: Here, we summarize the relations between our criteria and all these mentioned in the main text after applying LFT.
+After applying LFT, for the optimal parameter values, 1 coincides with Sarbicki et al.’s criterion, i.e.
+$\|\widetilde{\mathcal{T}}\|_{\mathrm{tr}}\leq 2\sqrt{(d_{A}-1)(d_{B}-1)/d_{A}d_{B}}$, which reproduces the Proposition 6 of Ref.
+[26] when $d_{A}=d_{B}$ and de Vicente’s criterion.
+The ESIC criterion is stronger than CCNR and they coincide as $d_{A}=d_{B}$ after applying LFT.
+Criteria $M_{A}/M_{B}$ $\mathcal{A}/\mathcal{B}$ $\kappa_{A}\kappa_{B}-\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}$
+de Vicente’s criterion $\left[\begin{smallmatrix}0&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}$
+Sarbicki et al.’s criterion $\left[\begin{smallmatrix}h&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4(d_{A}-1+h_{A}^{2})(d_{B}-1+h_{B}^{2})}{d_{A}d_{B}}}-\frac{|h_{A}||h_{B}|}{d_{A}d_{B}}$
+This work $\left[\begin{smallmatrix}\frac{t}{\sqrt{2d}}&\cdots&\frac{t}{\sqrt{2d}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}-|t_{A}||t_{B}|$
+CCNR $\left[\begin{smallmatrix}\frac{1}{\sqrt{d^{2}-1}}&\cdots&\frac{1}{\sqrt{d^{2}-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\frac{2(d_{A}d_{B}-\sqrt{d_{A}d_{B}})}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
+ESIC $\left[\begin{smallmatrix}\frac{1}{\sqrt{d-1}}&\cdots&\frac{1}{\sqrt{d-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\frac{4d_{A}d_{B}-2\sqrt{d_{A}d_{B}(d_{A}+1)(d_{B}+1)}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
+Proposition 6 in [26] $\left[\begin{smallmatrix}0&\cdots&0\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4d_{A}d_{B}}{(d_{A}+1)(d_{B}+1)}}$
+6 Relations to entanglement witness
 An entanglement witness is a Hermitian operator which has nonnegative expectation values for all separable states and negative one for some entangled states [7, 63].
 Entanglement witness provides an experimental friendly tool to analyse entanglement [3].
-To this end, we reformulate the Theorem 1 in terms of entanglement witness.
-Given a correlation matrix $\mathcal{C}$ for measurement vectors $\boldsymbol{A},\boldsymbol{B}$, we readily find
-$\|\mathcal{C}\|_{\mathrm{tr}}=\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\;.$ (29)
-Here, $A^{\prime}_{\mu}=\sum_{\nu}P_{\nu\mu}A_{\nu}$, $B^{\prime}_{\mu}=\sum_{\nu}Q_{\nu\mu}B_{\nu}$ and $P,Q$ achieve a singular value decomposition of $\mathcal{C}$, i.e.
-$\mathcal{C}=P\Sigma Q^{\mathrm{T}}$, where $P,Q$ are orthogonal matrices, $\Sigma=\operatorname{diag}\{\sigma_{1},\cdots,\sigma_{r},0,\cdots,0\}$ with singular values $\sigma_{k}$ and $\operatorname{rank}(\mathcal{C})=r$ [54].
+We next reformulate the Theorem 1 in terms of entanglement witness.
+To this end, we provide a new understanding for the trace norm of correlation matrix for observables $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$.
+Similar to the orbit of a set under group action [64], one can define the orbit of $m$-tuple $\boldsymbol{A}$ under the orthogonal group $O(m)$
+$\mathcal{O}(\boldsymbol{A})=\{O\boldsymbol{A}|O\in O(m)\}\;.$ (25)
+Here, $\boldsymbol{A}^{\prime}\in\mathcal{O}(\boldsymbol{A})$ implies $A^{\prime}_{\mu}=\sum_{\nu}O_{\mu\nu}A_{\nu},\mu=1,\cdots,m$.
+It is readily to prove that the all elements of $\mathcal{O}(\boldsymbol{A})$ yield the same MIBS up to a rotation.
+We can define the orbit $\mathcal{O}(\boldsymbol{B})$ of $n$-tuple $\boldsymbol{B}$ as well.
+Naturally, given $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$, the orbit of the correlation matrix is defined as
+$\mathcal{O}(\mathcal{C})\!\equiv\!\{\braket{\!A_{\mu}\otimes B_{\nu}\!}|\boldsymbol{A}\!\in\!\mathcal{O}(\boldsymbol{A}),\boldsymbol{B}\!\in\!\mathcal{O}(\boldsymbol{B})\!\}\;.$ (26)
+Without loss of generality, assume $m\geq n$.
+By padding $\boldsymbol{B}$ with zero matrices, we can reshape it into an $m$-tuple.
+Consequently, the trace norm of the correlation matrix can be reformulated as follows (see Appendix G for proof)
+$\|\mathcal{C}\|_{\mathrm{tr}}=\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]\;.$ (27)
+The von Neumann’s trace theorem [65] tells us that the maximum is reached when $\boldsymbol{A}^{\prime}=P\boldsymbol{A}$ and $\boldsymbol{B}^{\prime}=Q\boldsymbol{B}$, where $P,Q$ achieve a singular value decomposition (SVD) of $\mathcal{C}$, i.e.
+$\mathcal{C}=P\Sigma Q^{\mathrm{T}}$.
+Then, we have
+$\|\mathcal{C}\|_{\mathrm{tr}}=\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]=\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\;.$ (28)
+Here, $r=\operatorname{rank}(\mathcal{C})$.
+Eq.
+28 shows that there exist the optimal observables $A^{\prime}_{\mu}$ and $B^{\prime}_{\mu}$ to witness entanglement.
 Thus, one may immediately get an inequality of correlation function similar to CHSH inequality Eq.
@@
 $\mathcal{W}=\kappa\mathds{1}\otimes\mathds{1}-\sum_{\mu=1}^{r}A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}\;.$ (30)
-This means that based upon the extended Bloch space, entanglement witness will be simply generated for generic observables.
-Further extensions
-Through the construction of a measurement-induced Bloch space, we have established a general condition for separability which arises from the constraint of the length of the MIBVs.
-However, not only the moduli of MIBVs but also their directions suffice a constraint.
-As an illustration, we consider the measurement vector $\boldsymbol{A}=\boldsymbol{B}=\vec{\pi}$, in which case the measurement-induced Bloch space reduces to the familiar Bloch space.
-Then the angel $\theta$ between any two Bloch vectors of pure states satisfies $-\frac{1}{d-1}\leq\cos\theta\leq 1$ [55, 35].
-Given a $d\times d$ bipartite state, if it is separable, i.e.
-$\mathcal{T}=\sum_{i}p_{i}\vec{r}_{i}\vec{s}_{i}^{\mathrm{T}}$ with $\vec{r}_{i}$ and $\vec{s}_{i}$ being the Bloch vectors, we find the criterion
-$-\frac{2}{d}\leq\mathrm{Tr}[\mathcal{T}]\leq\frac{2(d-1)}{d}$ (32)
-exists, otherwise the entanglement survives.
-We notice $\mathrm{Tr}[\mathcal{T}]=\sum_{\mu=1}^{d^{2}-1}\braket{\pi_{\mu}\otimes\pi_{\mu}}$, so Eq.
-32 yields the entanglement witness $\mathcal{W}=\frac{2}{d}\mathds{1}\otimes\mathds{1}+\sum_{\mu}\pi_{\mu}\otimes\pi_{\mu}$.
-Considering arbitrary dimensional Werner states [47, 56]
-$\rho_{\mathrm{W}}=\frac{1}{d^{2}}\mathds{1}\otimes\mathds{1}+\frac{1}{4}\sum_{\mu=1}^{d^{2}-1}\frac{2(d\phi-1)}{d(d^{2}-1)}\pi_{\mu}\otimes\pi_{\mu}\;,\;\forall\phi\in[-1,1]\;,$
-which is separable if and only if $\phi\geq 0$ [47].
-An simple calculation shows that Eq.
-32 is a necessary and sufficient condition taking effect in full parameter space $\phi\in[-1,0)$ , while CCNR and ESIC criteria are valid merely in $\phi\in[-1,-\frac{d-2}{d}]$.
-With the growth of dimensionality, the entanglement detectable intervals of CCNR and ESIC criteria tend to none.
-This example tells that we can establish more powerful separability condition than existing ones by only employing the inherent constraints in Bloch space.
-On this account, we are tempting to conclude that the inherent constraints in Bloch space may address a lot of the separability issues.
-Concluding remarks
+This implies that, based on the extended Bloch space, entanglement witnesses can be easily generated for generic observables.
+7 Concluding remarks
 In the framework of measurement-induced Bloch space, by dint of a novel scheme we obtain a family of stronger separability criteria for entangled system.
@@
 Acknowledgements
-This work was supported in part by the National Natural Science Foundation of China(NSFC) under the Grants 11975236, 12235008, the Fundamental Research Funds for the Central Universities and by the University of Chinese Academy of Sciences.
-Supplemental Material
-Here we elaborate and derive in detail the main results discussed in the main text.
-The proof of Theorem 1
-Given arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, the correlation matrices of the separable states can be expressed as
+This work was supported in part by the National Natural Science Foundation of China(NSFC) under the Grants 12475087, 12235008, the Fundamental Research Funds for the Central Universities and by the University of Chinese Academy of Sciences, and China Postdoctoral Science Foundation funded project No.
+2024M753174.
+Appendix A The proof of Theorem 1
+Given $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, the correlation matrices of the separable states can be expressed as
 $\mathcal{C}=\sum_{i}p_{i}(\vec{t}\,^{A}/d_{A}+\vec{\alpha}_{i})(\vec{t}\,^{B}/d_{B}+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (31)
-Here, $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are the generalized Bloch vectors associated with $\boldsymbol{A}$ and $\boldsymbol{B}$ as defined in the main text.
+Here, $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are the measurement-induced Bloch vectors (MIBVs) associated with $\boldsymbol{A}$ and $\boldsymbol{B}$ as defined in the main text.
 Employing the convexity property of trace norm $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$, we have
@@
 So, the correlation matrices of the separable states are limited by the size of the measurement-induced Bloch space (MIBS).
-Next, we prove that the MIBS is bounded and its boundary can be obtained simply.
-Given a measurement vector $\boldsymbol{A}$, quantum states can be transformed into the measurement-induced Bloch vector (MIBVs) which constitute the MIBS.
-It is found that the MIBVs satisfy the following equation [1]
-$\vec{\alpha}^{\mathrm{T}}\Omega^{-}\vec{\alpha}=\mathrm{Tr}[\rho^{2}]-\frac{1}{d}\;.$ (36)
-Here, $\Omega_{\mu\nu}\equiv\mathrm{Tr}[\bar{A}_{\mu}\bar{A}_{\nu}]=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$ and $\Omega^{-}$ denotes the Moore-Penrose inverse of $\Omega$.
-Eq.
-36 is an ellipsoid equation, which characterizes the MIBS, that is,
-$\mathcal{B}(\boldsymbol{A})=\left\{\vec{\alpha}\Big{|}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}^{\mathrm{T}}\Omega^{-}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}\leq 1\right\}\;.$ (37)
+And we have defined the parameter $\beta_{A}$
+$\beta_{A}=\sup_{\rho}\sum_{k=1}^{m}\braket{\bar{A}_{k}}_{\rho}^{2}=\sup_{\vec{\alpha}\in\boldsymbol{B}(A)}|\vec{\alpha}|^{2}$ (34)
+similarly for $\beta_{B}$, which yields
+$\|\mathcal{C}\|_{\operatorname{tr}}\leq\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\beta_{A}}\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\beta_{B}}\;.$ (35)
+Next, we show that how to obtain an upper bound of $\beta_{A/B}$.
+Given an $m$-tuple $\boldsymbol{A}$, quantum states can be transformed into the measurement-induced Bloch vector (MIBVs) which constitute the MIBS.
+It is proved that MIBS is a proper subset of the following hyper-ellipsoid
+$\mathcal{B}(\boldsymbol{A})\subseteq\left\{\vec{\alpha}\Big{|}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}^{\mathrm{T}}\Omega^{-}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}\leq 1\right\}\;.$ (36)
+Here, $\Omega^{-}$ denotes the Moore-Penrose inverse of $\Omega$ which is the unique matrix satisfying the following Moore-Penrose conditons [66]
+$\Omega\Omega^{-}\Omega=\Omega\;,\;\Omega^{-}\Omega\Omega^{-}=\Omega^{-}\;,$ (37)
+$(\Omega\Omega^{-})^{\dagger}=\Omega\Omega^{-}\;,\;(\Omega^{-}\Omega)^{\dagger}=\Omega^{-}\Omega\;.$ (38)
+The Moore-Penrose inverse $\Omega^{-}$ coincides with the matrix inverse when $\Omega$ is invertible.
 Matrix $\Omega$ completely depicts the properties of the ellipsoid defined by Eq.
-37, such as the lengths of the semi-axes are given by $\sqrt{\frac{(d\mathrm{Tr}[\rho^{2}]-1)\lambda_{i}}{d}}$, with $\lambda_{i}$ being the eigenvalues of $\Omega$; the eigenvectors of $\Omega$ determines the directions of the semi-axes [2].
+36, such as the lengths of the semi-axes are given by $\sqrt{\frac{(d\mathrm{Tr}[\rho^{2}]-1)\lambda_{i}}{d}}$, with $\lambda_{i}$ being the eigenvalues of $\Omega$; the eigenvectors of $\Omega$ determines the directions of the semi-axes [67].
 Obviously, Bloch space is a special case of $\boldsymbol{A}=\vec{\pi}$.
-So, for arbitrary measurement vector $\boldsymbol{A}$, we have
-$|\vec{\alpha}|^{2}\leq\frac{d-1}{d}\lambda_{\max}(\Omega)\;,$ (38)
-where $\lambda_{\max}(X)$ is the maximal eigenvalue of matrix $X$. Finally, we prove
-$\|\mathcal{C}\|_{\operatorname{tr}}\leq\kappa_{A}\kappa_{B}\;.$ (39)
-Here, $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, similarly for $\kappa_{B}$.
-The proof of Observation 1
+So, for arbitrary $\boldsymbol{A}$, we have
+$\beta_{A}\leq\frac{d-1}{d}\lambda_{\max}(\Omega)\;,$ (39)
+where $\lambda_{\max}(X)$ is the maximal eigenvalue of matrix $X$. In particular, we have $\beta_{A}=\frac{d-1}{d}\lambda_{\max}(\Omega)$, when Eq. 36 is a hypersphere.
+Appendix B The proof of Observation 1
 If $\boldsymbol{A}$ has the same trace and the traceless part corresponds to a regular simplex, namely
-$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (42)
+$M_{A}=\left[\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\;,$ (40)
 where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e.
-$\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\},\mu,\nu=1,\cdots,d_{A}^{2}$, then we have $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=2\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$, whose matrix form is
+$\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\},\mu,\nu=1,\cdots,d_{A}^{2}$.
+Then we have $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=2\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$, whose matrix form is
 $\Omega_{A}=\frac{2}{d_{A}^{2}-1}\left(d_{A}^{2}\mathds{1}-J_{d_{A}^{2}}\right)\;,$ (41)
@@
 $d_{A}^{2}$ and 00 ($d_{A}^{2}-1$-fold degeneracy).
-Hence, we have $\lambda_{\max}(\Omega_{A})=\frac{2d_{A}^{2}}{d_{A}^{2}-1}$, and
+So, the MIBS $\mathcal{B}(\boldsymbol{A})$ is a degenerative hypersphere with radius $\sqrt{\frac{2d_{A}}{d_{A}+1}}$.
+Hence, we have $\beta_{A}=\frac{2d_{A}}{d_{A}+1}$, and
 $\kappa_{A}=\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\;,$ (42)
 similarly for $\kappa_{B}$. So, for the separable state, we have
-$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;,$ (45)
+$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;.$ (43)
 Here, parameters $t_{A}$ and $t_{B}$ can be arbitrary real numbers.
@@
 $\rho=\frac{1}{4}\sum_{\mu,\nu}\chi_{\mu\nu}\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B}=\sum_{k}\sigma_{k}\widetilde{\Pi}_{k}^{A}\otimes\widetilde{\Pi}_{k}^{B}\;.$ (45)
-Here, $\sum_{k}\sigma_{k}=\frac{1}{2}\|\chi\|_{\mathrm{tr}}$ and $\{\widetilde{\Pi}_{k}^{A}\}$ and $\{\widetilde{\Pi}_{k}^{B}\}$ form orthonormal bases of the operator space.
+where $\sum_{k}\sigma_{k}=\frac{1}{2}\|\chi\|_{\mathrm{tr}}$.
+$\{\widetilde{\Pi}_{k}^{A}\}$ and $\{\widetilde{\Pi}_{k}^{B}\}$ form orthonormal bases of the operator space.
 CCNR criterion says that, if $\rho$ is separable, then we have $\sum_{k}\sigma_{k}\leq 1$ [14, 13], that is, $\|\chi\|_{\mathrm{tr}}\leq 2$.
 If $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ for $X=A,B$, then we have
-$M_{A}^{\mathrm{T}}M_{A}$ $=\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{1}\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{2}\\ \vdots&\vdots\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\frac{1}{\sqrt{d_{A}^{2}-1}}&\cdots&\frac{1}{\sqrt{d_{A}^{2}-1}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}\;.$ (54)
+$M_{A}^{\mathrm{T}}M_{A}$ $=\left[\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{1}\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{2}\\ \vdots&\vdots\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\left[\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\frac{1}{\sqrt{d_{A}^{2}-1}}&\cdots&\frac{1}{\sqrt{d_{A}^{2}-1}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}\;.$ (46)
 Here, $\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$ has been employed.
@@
 which reproduces the ESIC criterion due to $\braket{A_{\mu}\otimes B_{\nu}}=\frac{4d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}\braket{E_{\mu}^{A}\otimes E_{\nu}^{B}}$.
-The vertex coordinates for regular simplex
+Appendix C The vertex coordinates for regular simplex
 There are many methods to generate the vertex coordinates for regular simplex, where we give a simple scheme for the reader’s convenience. Given the following $n$-dimensional unit vectors
@@
 then $\{\vec{e}_{1},\vec{e}_{2}\cdots,\vec{e}_{n+1}\}$ constitutes a $n$-dimensional regular simplex. The following is an example for $n=8$,
-$\left(\begin{array}[]{ccccccccc}1&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}\\ 0&\frac{3\sqrt{7}}{8}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}\\ 0&0&\frac{3}{2}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}\\ 0&0&0&\frac{\sqrt{15}}{4}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}\\ 0&0&0&0&\frac{3}{\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}\\ 0&0&0&0&0&\frac{3}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}\\ 0&0&0&0&0&0&\frac{\sqrt{3}}{2}&-\frac{\sqrt{3}}{4}&-\frac{\sqrt{3}}{4}\\ 0&0&0&0&0&0&0&\frac{3}{4}&-\frac{3}{4}\\ \end{array}\right)\;.$ (67)
-Bound entangled states
+$\left[\begin{array}[]{ccccccccc}1&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}\\ 0&\frac{3\sqrt{7}}{8}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}\\ 0&0&\frac{3}{2}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}\\ 0&0&0&\frac{\sqrt{15}}{4}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}\\ 0&0&0&0&\frac{3}{\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}\\ 0&0&0&0&0&\frac{3}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}\\ 0&0&0&0&0&0&\frac{\sqrt{3}}{2}&-\frac{\sqrt{3}}{4}&-\frac{\sqrt{3}}{4}\\ 0&0&0&0&0&0&0&\frac{3}{4}&-\frac{3}{4}\\ \end{array}\right]\;.$ (59)
+Appendix D The proof of Observation 2
+Via 1, one immediately obtain the first inequality for the $d\times d$ system, i.e.
+$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;.$ (60)
+In the main text, we reformulate the correlation matrix of the separable state based on the MIBVs, i.e.
+Eq.
+19.
+When the involved two subsystems share the same dimension and the traceless parts of $\boldsymbol{A}=\boldsymbol{B}$ correspond to a regular simplex with trace parts $t_{\mu}^{A}=t_{\mu}^{B}=t$, we have
+$\mathcal{C}=\sum_{i}q_{i}(\vec{t}/d+\vec{\alpha}_{i})(\vec{t}/d+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (61)
+Here, $\vec{t}=(t,t,\cdots,t)^{\mathrm{T}}$ and $\vec{\alpha}_{i},\vec{\beta}_{i}\in\mathcal{B}(\boldsymbol{A})$ with $\alpha_{i\mu}=\vec{r}_{i}\cdot\vec{a}_{\mu},\beta_{i\mu}=\vec{s}_{i}\cdot\vec{a}_{\mu}$, where $\vec{r}_{i}$ and $\vec{s}_{i}$ are Bloch vectors.
+And then, we have
+$\mathrm{Tr}[\mathcal{C}]$ $=\sum_{i}q_{i}\left[t^{2}+\frac{\vec{t}\cdot\vec{\alpha}_{i}+\vec{t}\cdot\vec{\beta}_{i}}{d}+\vec{\alpha}_{i}\cdot\vec{\beta}_{i}\right]$ (62)
+$=t^{2}+\sum_{i}q_{i}\vec{\alpha}_{i}\cdot\vec{\beta}_{i}\;,\;(\vec{t}\cdot\vec{\alpha}_{i}=\vec{t}\cdot\vec{\beta}_{i}=0)$ (63)
+$=t^{2}+\sum_{i}q_{i}(\vec{r}_{i}\cdot\vec{a}_{1},\vec{r}_{i}\cdot\vec{a}_{2},\cdots,\vec{r}_{i}\cdot\vec{a}_{d^{2}})\left(\begin{matrix}\vec{s}_{i}\cdot\vec{a}_{1}\\ \vec{s}_{i}\cdot\vec{a}_{2}\\ \vdots\\ \vec{s}_{i}\cdot\vec{a}_{d^{2}}\end{matrix}\right)$ (64)
+$=t^{2}+\sum_{i}q_{i}\left[\left(\begin{matrix}\vec{a}_{1}^{\mathrm{T}}\\ \vec{a}_{2}^{\mathrm{T}}\\ \vdots\\ \vec{a}_{d^{2}}^{\mathrm{T}}\end{matrix}\right)\vec{r}_{i}\right]^{\mathrm{T}}\left[\left(\begin{matrix}\vec{a}_{1}^{\mathrm{T}}\\ \vec{a}_{2}^{\mathrm{T}}\\ \vdots\\ \vec{a}_{d^{2}}^{\mathrm{T}}\end{matrix}\right)\vec{s}_{i}\right]$ (65)
+$=t^{2}+\sum_{i}q_{i}\vec{r}_{i}^{\mathrm{T}}\mathcal{A}\vec{s}_{i}\;,\;(\mathcal{A}=\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}=\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1})$ (66)
+$=t^{2}+\frac{d^{2}}{d^{2}-1}\sum_{i}q_{i}\vec{r}_{i}\cdot\vec{s}_{i}\;.$ (67)
+Due to $0\leq\mathrm{Tr}[\rho\rho^{\prime}]\leq 1$ with $\rho=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{r}_{i}\cdot\vec{\pi}$ and $\rho^{\prime}=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{s}_{i}\cdot\vec{\pi}$, one can readily find $-\frac{2}{d}\leq\vec{r}_{i}\cdot\vec{s}_{i}\leq\frac{2(d-1)}{d}$ which results in
+$t^{2}-\frac{2d}{d^{2}-1}\leq\mathrm{Tr}[\mathcal{C}]\leq t^{2}+\frac{2d}{d+1}\;.$ (68)
+Because of $\mathrm{Tr}[\mathcal{C}]\leq\|\mathcal{C}\|_{\mathrm{tr}}$, we only need to consider the left hand side of this inequality.
+In all, we have proved that if a $d\times d$ bipartite state is separable, then it satisfies
+$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;,$ (69)
+and
+$\mathrm{Tr}[\mathcal{C}]\geq t^{2}-\frac{2d}{d^{2}-1}\;.$ (70)
+Appendix E Bound entangled states
 Horodecki’s bound entangled states with white noise:
@@
 Here, $\rho_{\mathrm{H}}(t)$ is $3\times 3$ Horodecki’s bound entangled states [68], i.e.
-$\rho_{\mathrm{H}}(s)=\frac{1}{8s+1}\left(\begin{matrix}s&0&0&0&s&0&0&0&s\\ 0&s&0&0&0&0&0&0&0\\ 0&0&s&0&0&0&0&0&0\\ 0&0&0&s&0&0&0&0&0\\ s&0&0&0&s&0&0&0&s\\ 0&0&0&0&0&s&0&0&0\\ 0&0&0&0&0&0&\frac{1}{2}(1+s)&0&\frac{1}{2}\sqrt{1-s^{2}}\\ 0&0&0&0&0&0&0&s&0\\ s&0&0&0&s&0&\frac{1}{2}\sqrt{1-s^{2}}&0&\frac{1}{2}(1+s)\end{matrix}\right)\;.$ (78)
+$\rho_{\mathrm{H}}(s)=\frac{1}{8s+1}\left[\begin{matrix}s&0&0&0&s&0&0&0&s\\ 0&s&0&0&0&0&0&0&0\\ 0&0&s&0&0&0&0&0&0\\ 0&0&0&s&0&0&0&0&0\\ s&0&0&0&s&0&0&0&s\\ 0&0&0&0&0&s&0&0&0\\ 0&0&0&0&0&0&\frac{1}{2}(1+s)&0&\frac{1}{2}\sqrt{1-s^{2}}\\ 0&0&0&0&0&0&0&s&0\\ s&0&0&0&s&0&\frac{1}{2}\sqrt{1-s^{2}}&0&\frac{1}{2}(1+s)\end{matrix}\right]\;.$ (72)
 Unextendible product bases (UPB) bound entangled states:
@@
 In the main text, we test Observation 1 by randomly generated $50000$ chessboard states, where the six parameters are drawn independently from the standard normal distribution.
-The proof of Theorem 2
+Appendix F The proof of Theorem 2
 By virtue of Bloch representation, the correlation matrix can be written as $\mathcal{C}=M_{A}^{\mathrm{T}}\chi M_{B}$, so we have
@@
 Given arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, we have
-$M_{A}=\left(\begin{matrix}\frac{t_{1}^{A}}{\sqrt{2d_{A}}}&\frac{t_{2}^{A}}{\sqrt{2d_{A}}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;,\;M_{B}=\left(\begin{matrix}\frac{t_{1}^{B}}{\sqrt{2d_{B}}}&\frac{t_{2}^{B}}{\sqrt{2d_{B}}}&\cdots\\ \vec{b}_{1}&\vec{b}_{2}&\cdots\end{matrix}\right)\;,$ (89)
+$M_{A}=\left[\begin{matrix}\frac{t_{1}^{A}}{\sqrt{2d_{A}}}&\frac{t_{2}^{A}}{\sqrt{2d_{A}}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right]\;,\;M_{B}=\left[\begin{matrix}\frac{t_{1}^{B}}{\sqrt{2d_{B}}}&\frac{t_{2}^{B}}{\sqrt{2d_{B}}}&\cdots\\ \vec{b}_{1}&\vec{b}_{2}&\cdots\end{matrix}\right]\;,$ (79)
 and
-$M_{A}M_{A}^{\mathrm{T}}=\left(\begin{matrix}\frac{(\vec{t}^{A})^{2}}{2d_{A}}&0\\ 0&\mathcal{A}\end{matrix}\right)=\frac{(\vec{t}^{A})^{2}}{2d_{A}}\oplus\mathcal{A}\;,$ (92)
+$M_{A}M_{A}^{\mathrm{T}}=\left[\begin{matrix}\frac{(\vec{t}^{A})^{2}}{2d_{A}}&0\\ 0&\mathcal{A}\end{matrix}\right]=\frac{(\vec{t}^{A})^{2}}{2d_{A}}\oplus\mathcal{A}\;,$ (80)
 where $\mathcal{A}=\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}$.
@@
 $\mathrm{Tr}\left[\sqrt{\mathcal{A}\widetilde{\mathcal{T}}\mathcal{B}\widetilde{\mathcal{T}}^{\mathrm{T}}}\right]\leq\kappa_{A}\kappa_{B}-\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}\;.$ (84)
+Appendix G The proof of Eq. 28
+Without loss of generality, assume $m\geq n$. We add some zero matrices to reshape $\boldsymbol{B}$ into $m$-tuple, then we have
+$\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]=$ $\max_{\{O^{A},O^{B}\}}\sum_{\mu}\Braket{\sum_{\mu^{\prime}}O_{\mu\mu^{\prime}}^{A}A_{\mu^{\prime}}\otimes\sum_{\nu^{\prime}}O_{\mu\nu^{\prime}}^{B}B_{\nu^{\prime}}}$ (85)
+$=\max_{\{O^{A},O^{B}\}}\sum_{\mu}\sum_{\mu^{\prime}\nu^{\prime}}O_{\mu\mu^{\prime}}^{A}O_{\mu\nu^{\prime}}^{B}\braket{A_{\mu^{\prime}}\otimes B_{\nu^{\prime}}}$ (86)
+$=\max_{\{O^{A},O^{B}\}}\sum_{\mu}\sum_{\mu^{\prime}\nu^{\prime}}O_{\mu\mu^{\prime}}^{A}\mathcal{C}_{\mu^{\prime}\nu^{\prime}}(O^{B})^{\mathrm{T}}_{\nu^{\prime}\mu}$ (87)
+$=\max_{\{O^{A},O^{B}\}}\mathrm{Tr}[O^{A}\mathcal{C}(O^{B})^{\mathrm{T}}]$ (88)
+$=\|\mathcal{C}\|_{\mathrm{tr}}\;.$ (89)
+In the last line, we have employed von Neumann’s trace theorem [65].